Circular No. 6962

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1621–47 S. Dieters and P. Woods, University of Alabama, Huntsville (UAH); C.

Kouveliotou, Universities Space Research Association; and J. van Paradijs, UAH and University of Amsterdam, report: “We have analyzed 9200 s of RXTE pointed observations made on June 26 of SGR 1627–41 (IAUC 6944), which were centered on the possibly-associated supernova remnant G337.0–0.1 (cf. IAUC 6948, 6950). We detected one burst, confirming flux detection from the new source and strengthening the association of the two sources. Power spectral analysis of the observations reveals a quasi-periodic oscillation (QPO) peak centered at 0.15 Hz (∼ 6.7 s), with an r.m.s. of at least 3.2 percent, depending upon background and source contributions to the signal. Also within the RXTE field-of-view is the black-hole candidate 4U 1630–47, from which QPOs have been observed (IAUC 6823). However, no QPOs near the above frequency have been reported, and the intensity of 1630–47 was low during our observations. Further observations are required to determine which source is the origin of the QPO.”

COMET C/1998 K5 (LINEAR) D. H. Wooden, Ames Research Center, NASA; D. E. Harker and C. E.

Woodward, University of Wyoming; and W. L. Raas, SETI Institute and University of California at Los Angeles, report 10-μm spectrophotometry (3′′ aperture, resolution ∼ 220) of comet C/1998 K5, obtained with the Wyoming Infrared Observatory’s 2.34-m telescope (+ NASA/Ames HIFOGS) during the period June 28–30 UT. No silicate feature (9.0–11.6 μm) was detected in excess of the continuum, well-fitted by a blackbody with a color temperature 310 ± 10 K and a continuum flux density at 7.8 μm of 7.2 ± 0.4 × 10^{−18} W cm^{−2} μm^{−1}. The peak flux density measured at 10 μm was 7.5 ± 1.0 × 10^{−18} W cm^{−2} μm^{−1}.

Ephemeris extension to IAUC 6940 (elements MPEC 1998-M30):

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 6 0^{h}38.^{m}97 +51^{o}53.^{′}6 0.247 0.984 75.^{o}4 90.^{o}5 12.9
11 1 26.16 +49 00.4 0.280 0.970 72.5 91.5 13.1
16 1 59.99 +46 01.4 0.316 0.964 71.5 90.4 13.3
21 2 25.24 +43 16.6 0.352 0.966 71.7 88.0 13.6
26 2 44.89 +40 49.9 0.388 0.975 72.9 84.7 13.8
31 3 00.68 +38 40.2 0.424 0.992 74.8 80.9 14.1
Aug. 5 3 13.63 +36 44.8 0.458 1.016 77.1 76.8 14.4
1998 July 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green